# Long-Range Interacting Particles on a Helix: A Statistical and Correlation Analysis of Equilibrium Configurations

### v1

@@ IV.1.2 Cluster correlations
The clusterization of $\Phi_{0}^{i}$ can be further used as a metric to understand the inherent structure of sequences $\mathbf{\Phi}_{0}=(\Phi_{0}^{1},\Phi_{0}^{2},\ldots,\Phi_{0}^{N-1})$. Therefore, Fig. 5(a) shows the probability on how many distinct clusters the $\Phi_{0}^{i}$ of an EC are distributed (Clusters/EC or number of clusters per EC) for $N$ in the range of 2-8. Evidently, the number of internal coordinates per EC ($N-1$) serve as an upper limit for the number of clusters per EC. This results in a non-zero probability of an incremental number of clusters per EC with each increase in $N$ for $2\leq N\leq 6$. However, these probabilities are decreasing as $N$ increases. This is due to the instability of configurations with large interparticle distances as $N$ increases, as well as the simultaneous increase in configurations with low interparticle distances. This ultimately results in clusters whose widths are uneven, thereby allowing for fewer combinations where all $\Phi_{0}^{i}$ are in different clusters within an EC. Finally, for $N>6$ this behavior is interrupted as for $N=7$, the occupation of 5 clusters per EC would be theoretically possible but does not occur. For $N=8$ the maximum occuring number of Clusters/EC is then equal to the total number clusters (see Fig. 4(a-g)). Apart from the trivial two-body case, a high probability for two clusters per EC is observed for all $N$. For $N=3$ a similarly high probability occurs for one cluster per EC while for $4\leq N\leq 6$, there is a higher probability of more than two clusters per EC than of just one. The opposite is true for $N=7,8$, which simply follows from the large number of low interparticle distances that occupy the first and second clusters (see Fig. 4).
@@
To gain full statistical insight into the ordering of interparticle distances within an EC, Fig. 7(a-f) further provides positional information, by showing the probability of the cluster value $C(\Phi_{0}^{i})$, contingent on the index $i$, i.e., the $i$-th coordinate of an EC. A general trend that emerges from the transition from $N=3$ to $N=8$, is the shift in the probability distributions for $C(\Phi_{0}^{i})$ from nearly even to highly uneven distributions. However, the distribution for the first coordinate remains relatively even across all values of $N$. In contrast, the illustration reveals a consistently high and increasing probability that the $(N-1)$-th distance falls within the first cluster as $N$ increases. However, a more intriguing observation is that, with an increase in $N$, the probability that the interparticle distances between the inner particles ($i=2,\ldots,N-2$) exhibit a stronger increasing probability of being in cluster 1. Concurrently, for $4\leq N\leq 8$, the highest occupiable cluster is exclusively reached by the first ($i=1$) and the last coordinate ($i=N-1$). Therefore, stable ECs occur with a high probability in circumstances where the inner particles are in closer proximity to one another than the outermost particles are to their neighbored particle. Concomitantly, Fig. 7(a-f) reveals a weak symmetry between the probabilities of $C(\Phi_{0}^{i})$ and $C(\Phi_{0}^{N-i})$, indicating an inherent symmetry between the values of $\Phi_{0}^{i}$ and $\Phi_{0}^{N-i}$ in at least some ECs. Separately, an overall high probability for a hierarchy of the values $\Phi_{0}^{i}$ within the sequence of an EC from high interparticle distances for $i=1$ to smaller ones for $i=N-1$, is evident. In summary, there is a high probability for ECs to be either symmetric in $\Phi_{0}^{i}$ and $\Phi_{0}^{N-i}$ or hierarchically ordered in $\Phi_{0}^{i}$.
IV.1.3 Euclidean distances of ECs in the configuration space
@@ IV.2 Energetical properties of equilibria
The subsequent part will undertake a detailed examination of the potential energies associated with the ECs that have been previously analyzed. Fig. 9(a) illustrates the distribution of potential energies $E\equiv V$ for $N=2-8$. Each illustrated line corresponds to the energy of an EC and is referred to as an energy level. In accordance with the cumulative potential energy term a trend of increasing energy levels as $N$ increases can be identified. Nonetheless, the overlap of energy ranges for systems with different N, as well as the total energy range, increases with increasing $N$. However, the number of ECs is growing faster than the total energy range, consequently leading to an overall decrease in the mean spacing $\langle\Delta E\rangle$ of the energy levels. This is shown in the upper inset. Contrary to the prevailing trend, an increase in the mean spacing is observed in the transitions from $N=2$ to $N=3$ and from $N=4$ to $N=5$. Especially the former transition differs significantly; however, it represents an outlier in the statistical analysis of the few-body system.
@@ IV.3 Correlations between spatial and energetic properties
In order to expand the correlational analysis to the $(N-1)$-dimensional potential landscape, Fig. 11(a-g) shows the energy difference $\Delta E_{ij}=\lvert V(\mathbf{\Phi}_{0}^{i})-V(\mathbf{\Phi}_{0}^{j})\rvert$ in dependence on the pairwise distances $d_{ij}$ (see Fig. 8(a)). A correlation is manifested in the manner, that ECs that are in a large spatial distance also rather show a great difference in their energies, thereby underpinning the established energy hierarchy. In accordance with Fig. 9(a) and Fig. 8(a), the maximum energy level exhibits an upward trend, while the maximum pairwise distance demonstrates a downward trend as $N$ increases, resulting in a more and more pronounced increase in $\Delta E_{ij}$ by $d_{ij}$ for the crossover from $N=2$ to $N=8$.
@@ V Summary and Conclusions
A natural subsequent step would be to examine the dynamical properties of the helical few-body chain through a harmonic approximation. From a statistical viewpoint an extensive analysis of the phase space would be intriguing. Moreover, this could be combined with a more sophisticated optimization approach, utilizing the accumulated empiricism to study the many-body system.

### v2

@@ IV.1.2 Cluster correlations
The clusterization of $\Phi_{0}^{i}$ can be further used as a metric to understand the inherent structure of sequences $\mathbf{\Phi}_{0}=(\Phi_{0}^{1},\Phi_{0}^{2},\ldots,\Phi_{0}^{N-1})$. Therefore, Fig. 5(a) shows the probability on how many distinct clusters the $\Phi_{0}^{i}$ of an EC are distributed (Clusters/EC or number of clusters per EC) for $N$ in the range of 2-8. Evidently, the number of internal coordinates per EC (N1N-1) serve as an upper limit for the number of clusters per EC. This results in a non-zero probability of an incremental number of clusters per EC with each increase in $N$ for $2\leq N\leq 6$. However, these probabilities are decreasing as $N$ increases. This is due to the instability of configurations with large interparticle distances as $N$ increases, as well as the simultaneous increase in configurations with low interparticle distances. This ultimately results in clusters whose widths are uneven, thereby allowing for fewer combinations where all $\Phi_{0}^{i}$ are in different clusters within an EC. Finally, for $N>6$ this behavior is interrupted as for $N=7$, the occupation of 5 clusters per EC would be theoretically possible but does not occur. For $N=8$ the maximum occuring number of Clusters/EC is then equal to the total number clusters (see Fig. 4(a-g)). Apart from the trivial two-body case, a high probability for two clusters per EC is observed for all $N$. For $N=3$ a similarly high probability occurs for one cluster per EC while for $4\leq N\leq 6$, there is a higher probability of more than two clusters per EC than of just one. The opposite is true for $N=7,8$, which simply follows from the large number of low interparticle distances that occupy the first and second clusters (see Fig. 4).
@@
To gain full statistical insight into the ordering of interparticle distances within an EC, Fig. 7(a-f) further provides positional information, by showing the probability of the cluster value $C(\Phi_{0}^{i})$, contingent on the index $i$, i.e., the $i$-th coordinate of an EC. A general trend that emerges from the transition from $N=3$ to $N=8$, is the shift in the probability distributions for $C(\Phi_{0}^{i})$ from nearly even to highly uneven distributions. However, the distribution for the first coordinate remains relatively even across all values of $N$. In contrast, the illustration reveals a consistently high and increasing probability that the (N1)(N-1)-th distance falls within the first cluster as $N$ increases. However, a more intriguing observation is that, with an increase in $N$, the probability that the interparticle distances between the inner particles ($i=2,\ldots,N-2$) exhibit a stronger increasing probability of being in cluster 1. Concurrently, for $4\leq N\leq 8$, the highest occupiable cluster is exclusively reached by the first ($i=1$) and the last coordinate (i=N1i=N-1). Therefore, stable ECs occur with a high probability in circumstances where the inner particles are in closer proximity to one another than the outermost particles are to their neighbored particle. Concomitantly, Fig. 7(a-f) reveals a weak symmetry between the probabilities of $C(\Phi_{0}^{i})$ and $C(\Phi_{0}^{N-i})$, indicating an inherent symmetry between the values of $\Phi_{0}^{i}$ and $\Phi_{0}^{N-i}$ in at least some ECs. Separately, an overall high probability for a hierarchy of the values $\Phi_{0}^{i}$ within the sequence of an EC from high interparticle distances for $i=1$ to smaller ones for i=N1i=N-1, is evident. In summary, there is a high probability for ECs to be either symmetric in $\Phi_{0}^{i}$ and $\Phi_{0}^{N-i}$ or hierarchically ordered in $\Phi_{0}^{i}$.
IV.1.3 Helical configurations
Figure 8: Helical equilibrium configurations for $N=8$ with (a) $\mathbf{\Phi}_{0}=(1.64,1.72,1.73,1.73,1.73,1.72,1.64)$ and (b) $\mathbf{\Phi}_{0}=(0.26,0.25,0.23,0.17,0.23,0.25,0.26)$.
@@ IV.1.3 Helical configurations
In addition to the statistical analysis of interparticle distances, the occurrence of intrinsically helical ECs should be emphasized. This refers to ECs, in which the particle chain exhibits a helical configuration due to its inherent particle ordering. The assumption that an EC is helical is made under the following criteria for the $\Phi_{0}^{i}$ of an EC: $\Phi_{0}^{i}\neq 1$ and $\Phi_{0}^{1}\approx\Phi_{0}^{2}\approx\ldots\approx\Phi_{0}^{N-1}$, where deviations from exact equidistancy such that $\Phi_{0}^{i,\mathrm{max}}-\Phi_{0}^{i,\mathrm{min}}<0.2$ are tolerated. Therefore, these ECs can be roughly associated with the potential wells of the global diagonal of the potential landscape (see Fig. 2(b)). Such configurations are illustrated in Fig. 8 for $N=8$ and for $\Phi_{0}^{i}$ within the second (a) and within the first (b) cluster. The existence of stable ECs in the form of a helical geometry underscores its significance in the context of stable configurations of repulsive long-range interacting particle systems, not only in the case of helical confinement but also for analogous systems with cylindrical symmetric confinement (see Refs. [25, 26, 27]).
IV.1.4 Euclidean distances of ECs in the configuration space
@@ IV.2 Energetical properties of equilibria
The subsequent part will undertake a detailed examination of the potential energies associated with the ECs that have been previously analyzed. Fig. 10(a) illustrates the distribution of potential energies $E\equiv V$ for N=28N=2-8. Each illustrated line corresponds to the energy of an EC and is referred to as an energy level. In accordance with the cumulative potential energy term a trend of increasing energy levels as $N$ increases can be identified. Nonetheless, the overlap of energy ranges for systems with different N, as well as the total energy range, increases with increasing $N$. However, the number of ECs is growing faster than the total energy range, consequently leading to an overall decrease in the mean spacing $\langle\Delta E\rangle$ of the energy levels. This is shown in the upper inset. Contrary to the prevailing trend, an increase in the mean spacing is observed in the transitions from $N=2$ to $N=3$ and from $N=4$ to $N=5$. Especially the former transition differs significantly; however, it represents an outlier in the statistical analysis of the few-body system.
@@ IV.3 Correlations between spatial and energetic properties
In order to expand the correlational analysis to the (N1)(N-1)-dimensional potential landscape, Fig. 12(a-g) shows the energy difference $\Delta E_{ij}=\lvert V(\mathbf{\Phi}_{0}^{i})-V(\mathbf{\Phi}_{0}^{j})\rvert$ in dependence on the pairwise distances $d_{ij}$ (see Fig. 9(a)). A correlation is manifested in the manner, that ECs that are in a large spatial distance also rather show a great difference in their energies, thereby underpinning the established energy hierarchy. In accordance with Fig. 10(a) and Fig. 9(a), the maximum energy level exhibits an upward trend, while the maximum pairwise distance demonstrates a downward trend as $N$ increases, resulting in a more and more pronounced increase in $\Delta E_{ij}$ by $d_{ij}$ for the crossover from $N=2$ to $N=8$.
@@ V Summary and Conclusions
A natural subsequent step would be to examine the dynamical properties of the helical few-body chain through a harmonic approximation. From a statistical viewpoint an extensive analysis of the phase space would be intriguing. Moreover, this could be combined with a more sophisticated optimization approach, utilizing the accumulated empiricism to study the many-body system. In fact, incomplete simulations, extending to $N=20$, show the persistence of many of the features previously outlined, including the clustering of interparticle distances, the diminution of the width of higher clusters and the concomitant augmentation of the first cluster, which consequently predominates in the statistical and correlational analysis. Additional persistent features are for instance a high degree of irregularity within the spatial distribution of ECs as well as increasing energies. Nonetheless, it would be unwarranted to expect that the features of the system presented here would be equally present in systems with a significantly larger number of particles, which therefore could give rise to newly emerging features.
@@
Acknowledgements.
J.M.D. and P.S. appreciate very much helpful discussions with Ansgar Siemens in the early stages of this project.